show me how to solve what is 15% of 8/9?
Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

show me how to solve what is 15% of 8/9?
------------
Ans: (15/100)(8/9) = (5/100)(8/3) = (1/20)(8/3) = (1/5)(2/3) = 2/15
---------------
Cheers,
Stan H.
---------------

Answer by Boreal(15235)

One way is remember that 15%=15/100 or 3/20
3/20 * (8/9)=24/180
That is 2/15, dividing numerator and denominator by 12.